OpenStudy (anonymous):

critical points: -5y^2+6x-1=0 10y^2-10xy=0

OpenStudy (anonymous):

x= 1/5 y = .2 x = 1 y = 1

OpenStudy (anonymous):

how did you come on these values? could you please explain your steps?

OpenStudy (anonymous):

Differentiate both equations and equate them to 0you get -10y+6=0 and 20y-10y-10x=0 i.e., x=y and y=6/10=3/5 hence x=y=3/5 are the critical points.

OpenStudy (anonymous):

you know the basics of differentiation right?

OpenStudy (anonymous):

ofcourse i know the basics of differentiation.. but i havent done this... with the substitution